# Revisiting MAE pre-training for 3D medical image segmentation

### v1

@@ Abstract
Self-Supervised Learning (SSL) presents an exciting opportunity to unlock the potential of vast, untapped clinical datasets, for various downstream applications that suffer from the scarcity of labeled data. While SSL has revolutionized fields like natural language processing and computer vision, their adoption in 3D medical image computing has been limited by three key pitfalls: Small pre-training dataset sizes, architectures inadequate for 3D medical image analysis, and insufficient evaluation practices. We address these issues by i) leveraging a large-scale dataset of 44k 3D brain MRI volumes and ii) using a Residual Encoder U-Net architecture within the state-of-the-art nnU-Net framework. iii) A robust development framework, incorporating 5 development and 8 testing brain MRI segmentation datasets, allowed performance-driven design decisions to optimize the simple concept of Masked Auto Encoders (MAEs) for 3D CNNs. The resulting model not only surpasses previous SSL methods but also outperforms the strong nnU-Net baseline by an average of approximately 3 Dice points. Furthermore, our model demonstrates exceptional stability, achieving the highest average rank of 2 out of 7 methods, compared to the second-best method’s mean rank of 3. Our code is made available here.
@@ 1 Introduction
In recent years, the concept of Self-Supervised Learning (SSL) has emerged as a driving factor in data-rich domains, enabling large-scale pre-training that facilitates the learning of robust and transferable general-purpose representations (Assran et al., 2023; Oquab et al., 2023; He et al., 2022). This paradigm shift has been instrumental in advancing various fields, particularly in domains with abundant labeled data like NLP or natural vision. In the domain of 3D medical image computing, this trend has not caught on.
Currently, the domain is either focused on training from-scratch, mainly using the nnU-Net framework by Isensee et al. (2021), or using supervised pretraining, which is limited by the cost associated with annotated data (Wasserthal et al., 2023; Ulrich et al., 2023; Huang et al., 2023). The usage of supervised pretraining implies the willingness of the domain to adopt pretraining and calls to question the currently established SSL methods in the domain. We believe this lack of widespread adoption of previously established methods can be attributed to three major pitfalls of previous SSL research in this field:
@@
P1 - Limited dataset size: Many SSL approaches have been developed on very few unlabeled volumes, often being trained on fewer than 10,000 images Zhou et al. (2021); Wu et al. (2024); Tang et al. (2024; 2022); Zhuang et al. (2023); Hatamizadeh et al. (2021), almost approaching scales of supervised dataset sizes. These datasets tend to be pooled from publicly available annotated datasets, as larger datasets pose a greater hurdle to acquire. While many hospitals possess 3D medical images in the millions, they are locked away from the public due to patient privacy concerns. While some larger open-source datasets exist, e.g., the Adolescent Brain Cognitive Decline (ABCD) dataset of the NIH (N=40k) or the UK Biobank (UK-BB) (N=120k), they restrict access pending an internal review board’s approval, posing a hurdle for open science. More recently, the UK-BB stopped allowing downloading data to local hardware, signaling a public decrease in the community’s willingness to share data.
@@
P2 - Outdated Backbones: Many studies develop their SSL method on non-state-of-the-art architectures, e.g., utilizing transformers (Wu et al., 2024; Tang et al., 2022; Wang et al., 2023; Chen et al., 2023). While transformers are prevalent in the 2D natural imaging domain (Dosovitskiy, 2020), recent architectures leveraging attention (Vaswani, 2017) have so far not been able to reach state-of-the-art performance in 3D medical segmentation. In fact, well-configured 3D U-Net (Çiçek et al., 2016; Ronneberger et al., 2015) inspired CNNs dominate 3D medical image segmentation, outperforming transformer-based models by a large margin Isensee et al. (2024). This underscores the need for SSL methods that can be seamlessly integrated with CNNs to harness their full potential in medical image analysis on downstream tasks.
@@
P3 - Insufficient Evaluation: Existing methods often lack rigorous evaluation, masking the methods’ efficacy (or lack thereof). This is represented through: i) Evaluating on too few datasets to show generalization of the pre-training ii) Stacking multiple contributions, e.g., novel architecture with new pre-training, which does not allow one to draw conclusions if the pre-training is effective without the architecture (Wang et al., 2023) iii) Comparing to bad from-scratch baselines, like a badly configured, outdated Çiçek et al. (2016) model instead of a well-configured 3D nnU-Net CNN baseline (Isensee et al., 2021). iv) Evaluating their method on seen data they pre-trained on. We want to emphasize that we do not intend to point fingers, but to raise awareness that evaluation matters and insufficient evaluation can lead to a lack of clarity which methods are the best. This observation is similar to the recent study by Isensee et al. (2024), that shows that this is also prevalent in the field of medical image segmentation when training from-scratch.
@@
In this paper, we carefully avoided all these pitfalls while exploring the Masked Auto Encoder (MAE) paradigm for 3D CNNs, with the recent adaptations introduced by Tian et al. (2023); Woo et al. (2023), and highlight that MAEs are exceeding the current state-of-the-art SSL methods in 3D medical image segmentation given proper configuration. Our contributions can be summarized as follows:
@@
    We evade Pitfall 1 by leveraging a dataset collection of 44k 3D MRI volumes to develop our self-supervised pretraining, exceeding the majority of 3D medical image segmentation SSL methods in scale (Zhou et al., 2021; Wu et al., 2024; Tang et al., 2024; 2022; Zhuang et al., 2023; Hatamizadeh et al., 2021).
@@
    We evade Pitfall 2 by utilizing the state-of-the-art Residual Encoder U-Net CNN architecture from Isensee et al. (2024) as the backbone. Moreover, we use this backbone for all baseline SSL methods - allowing us to quantify the utility of the pre-training scheme for pre-training CNNs.
@@
    We evade Pitfall 3 by employing 5 development and 7 testing datasets, spanning a diverse set of downstream targets. This includes head and neck organs and pathologies, datasets with novel image modalities not seen during pre-training, and datasets of the same pathology acquired at different centers.
By subsequently optimizing MAEs for CNNs based on performance on the development pool, we propose the novel sparse MAE inspired pre-training paradigm, unleashing the potential of huge amounts of not annotated data in the medical domain. We show that the choice of the best fine-tuning strategy is crucial. Moreover, we evaluate our method across a range of critical 3D medical scenarios, including low-data regimes, accelerated fine-tuning schedules, and generalization across unseen centers and modalities.
Figure 1: Overview of the model development pipeline. During pretraining, we carefully develop our method with 44k images (P1), with a state-of-the-art residual encoder U-Net (P2) and evaluate it on dive development datasets (P3). In the testing phase we utilize a total of thirteen datasets (P3) to verify the utility of our developed method.
@@ 2 Development Framework
(i) The architecture used is always the same state-of-the-art residual encoder U-Net architecture (ResEnc U-Net) (Isensee et al., 2024). (ii) The input patch size is [160x160x160]. (iii) All images are resampled to the target spacing of [1x1x1] mm3 (Roy et al., 2023). (iv) All images are z-score normalized to zero mean and unit variance (Isensee et al., 2021). (v) As optimizer, we use SGD with a decreasing poly-learning rate (Chen et al., 2018) following nnU-Net (Isensee et al., 2021). (vi) We always employ random sampling during pre-training, irrespective of the prevalence of the different MR modalities.
@@
To develop our pre-training method, we utilize a proprietary brain MRI dataset sourced from over 44 centers containing over 9k patients comprising a total of approx. 44k 3D MRI scans. Due to the variety of data sources, this dataset contains images from more than 10 different MR scanners, various MR modalities and a diverse patient population. For more details on the distributions we refer to Figure 2.
@@
    MS FLAIR (D1): Consensus delineations of multiple sclerosis (MS) lesions on T2-weighted FLAIR images(Muslim et al., 2022).
@@
    Brain Mets (D2): Brain metastases imaged through T1, contrast enhanced gradient echo T1ce, contrast enhanced spin echo T1 and a T2 FLAIR sequence acquired in the Stanford University Hospital (Grøvik et al., 2020).
@@
    Hippocampus (D3): The Hippocampus dataset, task 4 of the medical segmentation decathlon (MSD) (Antonelli et al., 2022), contains delineations of the anterior and posterior Hippocampus in T1 weighted MRI (Simpson et al., 2019).
@@
    Atlas22 (D4): Anatomical Tracings of Lesions After Stroke (ATLAS) on T1 weighted images. We use the Atlas R2.0 dataset from Liew et al. (2022).
@@
    CrossModa (D5): Delineations of intra-meatal and extra-meatal vestibular schwannoma tumors and cochlea delineations in contrast enhanced T1 weighted MRI (Dorent et al., 2023).
Of all these datasets, we set aside a hold-out test set comprising 20% of all images before start of the method development process. The remaining images were further split 80/20 into training and validation sets for the development process.
@@
    Cosmos (D6): This dataset features carotid vessel wall segmentation and atherosclerosis diagnosis, of which we use the contours to evaluate segmentation performance (Chen et al., 2022).
@@
    HaNSeg (D7): This dataset contains segmentations of 30 organs at risk (OAR), with associated T1 MRI and CT, of which we only use the MR images for model development (Podobnik et al., 2023).
@@
    Isles22 (D8): This dataset contains annotations of ischemic stroke lesions, with associated diffusion-weighted imaging (DWI), apparent diffusion coefficient (ADC) and T2 FLAIR (Hernandez Petzsche et al., 2022).
@@
    HNTS-MRG24(D9): T2-weighted MR images of pre-treatment oropharyngeal cancer and metastatic lymph nodes with associated annotations (Wahid et al., 2024).
@@
For all test datasets that are fine-tuned on (D6-D11) we use an 80/20 split for fine-tuning and for testing, as we fine-tune each method only once without any interventions. For D13 all data is used, as no training is conducted.
[t]
Figure 2: Hierarchical composition of the nnSSL foundation model’s pre-training dataset. (a) Geographic distribution of 44 participating centers. (b) Patient demographics (age, sex, weight). (c) Scanner specifications (magnetic field strength, manufacturer, scanner) of examinations. (d) Acquisition parameters (orientation, contrast administration, weighting) of image volumes.
@@ 3 Revisiting 3D MAEs
(a) Sparsification: Introducing all sparsification adaptations showed best results, with most gains from adding the additional densification conv layer between encoder and decoder.
@@
(b) Masking ratio: Masking ratios between 60% and 75% worked best when choosing static ratios with a dynamic range including higher masking ratios performing the best.
Pretraining D1 D2 D3 D4 D5 Avg. D1-D5
No Dyn. 45.56 72.26 88.80 60.44 82.61 69.93
No Fixed 49.37 69.13 88.78 60.74 81.33 69.87
VoCo 50.35 67.20 88.22 57.82 80.29 68.77
VF 49.93 69.58 88.83 61.75 81.48 70.31
MG 50.50 71.14 88.83 63.29 82.15 71.18
S3D-B (ours) 51.49 74.01 88.83 62.39 81.54 71.65
S3D-L (ours) 51.42 72.84 89.09 63.30 82.15 71.76
(c) Compound scaling: When comparing our scaled S3D-L, trained for more iterations with higher batch size and lr, only a modest increase in performance can be observed over the much more compute-efficient S3D-B. Moreover, when comparing models on the same scale S3D-B exceeds all baselines on our development validation datasets.
Masked autoencoders (MAEs) are a well-established pre-training paradigm in the natural imaging domain and in the medical image segmentation domain for transformers. In this section, we investigate this paradigm and optimize it for 3D medical image segmentation using a ResEnc U-Net architecture of Isensee et al. (2024).
@@
MAEs are trained by masking an input image to a certain degree and training the network to reconstruct the occluded regions, minimizing deviations betweem reconstruction and original image. In our experiments, we train the MAE with a L2-Loss in the z-score normalized voxel space and only calculate the reconstruction loss where regions were masked. Moreover, we do not remove skip-connections, following the general consensus of Woo et al. (2023), Tian et al. (2023), and He et al. (2022). The default hyperparameters (as used by the model denoted in gray in Table 1) are learning rate 1e-2, weight decay 3e-5, batch size 6, SGD optimizer with Nesterov momentum 0.99, masking ratio of 75% trained with a PolyLR schedule for 250k steps (this represent 1000 epochs in the nnU-Net framework) and minor spatial augmentations of affine scaling, rotation and mirroring.
@@
When masking the input image, CNNs are not able to ignore the masked regions in the same manner as transformers can. To address this, Tian et al. (2023) proposed to adapt the CNN architectures to better fit the sparse inputs: (a) Sparse Convolutions and Normalization:Through the receptive field of convolutions masked-out regions are iteratively eroded from their boundaries. By re-applying the masked regions after every convolution this problem can be resolved. Moreover, the masks can introduce a problematic shift in the normalization layer statistics due to the introduced zero values. To resolve this normalization is constrained to only consider the non-masked values. (b) Mask Token:Instead of feeding the light-weight decoder feature maps with zeroed mask regions, the regions are densified by filling them with a learnable mask token, simplifying the reconstruction task of the decoder. (c) Densification Convolution:After filling the masks with the Mask Token and before passing the feature maps to the decoder, a [3x3x3] convolution is applied to the feature maps at every resolution except the highest resolution to prepare the representations for decoding.
Results of these changes are visualized in Table 1(a). The adaptations are introduced iteratively, meaning the ’MaskToken’ ablation is only applied together with the Sparse Convolutions and Normalization. It can be observed that the full set of adaptations improves performance by an average of $0.3$ DSC points across our development datasets. Subsequently, all adaptations are kept and the following evaluations are presented with these changes applied.
@@
The masked region is determined by sampling randomly in the CNN’s bottleneck of shape [5x5x5] and up-sampling these regions to the input resolution, to ensure the masks align in the bottleneck of the CNN architectures. This results in masking regions of [32x32x32] voxels of non-overlapping regions in the input. As a sampling strategy, we follow random masking, as previous work showed no benefit of structured masking for images nor videos (He et al., 2022; Feichtenhofer et al., 2022). In the scope of the development phase we explore 5 static masking ratios between 30 and 90%, and evaluate a dynamic masking ratio randomly masking between 60% and 90%.
Results are presented in Table Table 1(b) and highlight that the masking ratios of 60%, 75%, and the dynamic masking ratio of 60% to 90% perform equally well. Due to the highly similar performance, we choose to proceed with a dynamic masking ratio over the static masking ratio, due to expecting this masking to be more difficult to learn and with the upcoming scaling experiment in mind. We refer to this model as Spark3D-Base (S3D-B= going forward.
Scaling
MAEs are known to benefit from scaling. We evaluate the effect of compound scaling by increasing the batch size by x8 to $48$, the learning rate to $23e-$ 2, and the iterations by x4 to $1M$. We refer to this model as S3D-Large (S3D-L) to denote the higher compute resources, but note that the architecture remains identical, to maintain easy adaptation of the parameters. Results are presented in Table 1(c). It can be observed that this x32 increase in compute only resulted in a slight increase in performance of 0.1 DSC points.
Fine-tuning strategy
Given a pre-trained model, a crucial question arises: Which weights to transfer and how to schedule the fine-tuning? We investigated various different schedules. Regarding weight transfer we investigated transferring (i) both the encoder and decoder [Uncaptioned image], or (ii) only the encoder [Uncaptioned image] with a randomly initialized decoder. Regarding the fine-tuning schedule, we investigate whether to use a learning rate warm-up of 12.5k steps, ramping the learning rate up to the maximum LR. When only transferring the encoder, an additional warm-up of only the decoder is investigated to adapt the randomly initialized decoder to the pre-trained encoder. In some configurations, this results in two learning rate warm-ups of 12.5k steps each. Additionally, we investigate whether to keep the encoder frozen [Uncaptioned image] for the entire fine-tuning process, or to fine-tuning encoder weights as well. Lastly, we investigate whether to decrease the peak learning rate to $31e-$ 3, $41e-$ 4 or to keep it at the default of $21e-$ 2.
Transfer 1. Warm-Up 2. Warm-Up Tr. Stage Max. LR D1 D2 D3 D4 D5 Avg

### v2

@@ Abstract
Self-Supervised Learning (SSL) presents an exciting opportunity to unlock the potential of vast, untapped clinical datasets, for various downstream applications that suffer from the scarcity of labeled data. While SSL has revolutionized fields like natural language processing and computer vision, its adoption in 3D medical image computing has been limited by three key pitfalls: Small pre-training dataset sizes, architectures inadequate for 3D medical image analysis, and insufficient evaluation practices. In this paper, we address these issues by i) leveraging a large-scale dataset of 39k 3D brain MRI volumes and ii) using a Residual Encoder U-Net architecture within the state-of-the-art nnU-Net framework. iii) A robust development framework, incorporating 5 development and 8 testing brain MRI segmentation datasets, allowed performance-driven design decisions to optimize the simple concept of Masked Auto Encoders (MAEs) for 3D CNNs. The resulting model not only surpasses previous SSL methods but also outperforms the strong nnU-Net baseline by an average of approximately 3 Dice points setting a new state-of-the-art. Our code and models are made available here.
@@ 1 Introduction
Figure 1: Well-configured MAE pre-training for CNNs is state-of-the-art. When comparing our MAE pre-trained model on all eleven test Datasets, Spark3D achieves almost +3 DSC points compared to the strong nnU-Net baseline and outperforms current SSL methods.
@@
In recent years, the concept of Self-Supervised Learning (SSL) has emerged as a driving factor in data-rich domains, enabling large-scale pre-training that facilitates the learning of robust and transferable general-purpose representations [2, 25, 14]. This paradigm shift has been instrumental in advancing various fields, particularly in domains with abundant labeled data like NLP or natural vision. In the domain of 3D medical image computing, this trend has not caught on. Currently, the domain is either focused on training from-scratch, mainly using the nnU-Net framework by Isensee et al. [18], or using supervised pre-training, which is limited by the cost associated with annotated data [40, 35, 17]. The usage of supervised pre-training implies the willingness of the domain to adopt pre-training and calls to question the currently established SSL methods in the domain. We believe this lack of widespread adoption of previously established methods can be attributed to three major pitfalls of previous SSL research in this field:
@@
P1 - Limited dataset size: Many SSL approaches have been developed on very few unlabeled volumes, often being trained on fewer than 10,000 images [44, 42, 30, 31, 45, 12], almost approaching scales of supervised dataset sizes. These datasets tend to be pooled from publicly available annotated datasets, as larger datasets pose a greater hurdle to acquire. While many hospitals possess 3D medical images in the millions, they are locked away from the public due to patient privacy concerns. Some larger open-source datasets exist, e.g., the Adolescent Brain Cognitive Decline (ABCD) dataset of the NIH (N=40k) or the UK Biobank (UK-BB) (N=120k), but they restrict access pending an internal review board’s approval, posing a hurdle for open science.
@@
P2 - Outdated Backbones: Many studies develop their SSL method on non-state-of-the-art architectures, e.g., utilizing transformers [42, 31, 39, 5]. While transformers are prevalent in the 2D natural imaging domain [8], recent architectures leveraging attention [36] have so far not been able to reach state-of-the-art performance in 3D medical segmentation. In fact, well-configured 3D U-Net [6, 27] inspired CNNs dominate 3D medical image segmentation, outperforming transformer-based models by a large margin [19]. This underscores the need for SSL methods that can be seamlessly integrated with CNNs to harness their full potential in medical image analysis on downstream tasks.
@@
P3 - Insufficient Evaluation: Existing methods often lack rigorous evaluation, masking the methods’ efficacy. This is represented through: i) Evaluating on too few datasets to show generalization of the pre-training ii) Stacking multiple contributions, e.g., novel architecture with new pre-training, which does not allow one to draw conclusions if the pre-training is effective without the architecture [39] iii) Comparing to bad from-scratch baselines, like a badly configured, outdated [6] model instead of a well-configured 3D nnU-Net CNN baseline [18]. iv) Evaluating their method on seen data they pre-trained on. We want to emphasize that we do not intend to point fingers, but to raise awareness that evaluation matters and insufficient evaluation can lead to a lack of clarity which methods are the best. This observation is similar to the recent study by Isensee et al. [19], that shows that this is also prevalent in the field of medical image segmentation when training from-scratch.
Table 1: Current SSL methods are limited by three pitfalls: They are trained on less than 10K 3D images (P1), they do not use a state-of-the-art backbone for 3D image segmentation as recommended by Isensee et al. [19] (P2) or they do not evaluate their method on more than 4 downstream datasets (P3). [Uncaptioned image] indicates evading a pitfall while [Uncaptioned image] indicates falling for a pitfall.
Method Year P1 P2 P3
Data Model Eval.
3DMAE [5] Jan. 2023 [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
GL-MAE [45] Aug. 2023 [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
MAPSeg [43] March 2023 [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
HySparK [30] Aug. 2024 [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
AMAES [23] Aug. 2024 [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
Swin UNETR [31] March 2022 [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
VoCo [42] April 2024 [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
VF [39] June 2023 [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
MG [44] Aug. 2019 [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
GVSL [15] March 2023 [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
S3D (ours) [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
@@
In this paper, we explore the Masked Auto Encoder (MAE) paradigm for 3D CNNs, with the recent adaptations introduced by Tian et al. [32], Woo et al. [41]. Our approach addresses the key pitfalls of previous SSL research in 3D medical image segmentation, ensuring broader applicability and higher performance. Our main contributions are as follows:
@@
    State-of-the-Art Performance: Our final model achieves superior performance compared to existing SSL-pre-trained models using the same architecture and dataset, see Fig. 1. Notably, this is the first work to demonstrate that SSL pretraining with a fixed architecture can consistently outperform a state-of-the-art, dynamically optimized nnU-Net baseline by approximately 3 Dice points across 11 diverse downstream datasets [18]. An associated model will be made available.
@@
    Comprehensive Evaluation Addressing Previous Pitfalls: Building on the detailed analysis of limitations in prior work (see Tab. 1), we set up a robust development environment to design our method. () P1: We utilize a collection of 39k 3D MRI volumes for self-supervised pre-training, substantially larger than most datasets used in prior SSL work [44, 42, 30, 31, 45, 12]. () P2: By using the state-of-the-art Residual Encoder U-Net CNN architecture [19], we bypass the limitations of outdated or transformer-based architectures, leveraging CNNs’ superior performance in 3D medical segmentation. () P3: We rigorously evaluate our model across five development and eight testing datasets, covering a wide range of downstream tasks, diverse pathologies, novel image modalities, and multi-center datasets. Moreover, we disentangle pre-training methodology from other contributions like novel backbones and include the strong 3D nnU-Net baseline, ensuring reliable results.
@@
    In-Depth Ablation Studies: We provide detailed analyses of critical design decisions, such as masking ratio, sparsification, and fine-tuning strategies. Additional ablations explore generalization, low-data performance, and faster fine-tuning, offering valuable insights.
While the fundamental concept of an MAE is not novel, its application in 3D medical image analysis has yet to be fully validated due to limitations in previous studies, which we illustrate in Tab. 1. Our carefully designed experiments close these gaps and reveal MAEs true potential.
@@ 2 Development Framework
(i) The architecture used is always the same state-of-the-art residual encoder U-Net architecture (ResEnc U-Net) [19, 35]. (ii) The input patch size is set to [160x160x160]. (iii) All images are resampled to the same target spacing of [1x1x1] mm [28]. (iv) All images are z-score normalized to zero mean and unit variance [18]. (v) As optimizer, we use SGD with a decreasing poly-learning rate [4] following nnU-Net [18]. (vi) We always employ random sampling during pre-training, irrespective of the prevalence of the different MR modalities.
@@
To develop our pre-training method, we utilize a proprietary brain MRI dataset sourced from over 44 centers containing over 9k patients comprising a total of approx. 44k 3D MRI scans. Due to the variety of data sources, this dataset contains images from more than 10 different MR scanners, various MR modalities and a diverse patient population. For more details on the distributions, we refer to Figure 3.
@@
    MS FLAIR (D1): Consensus delineations of multiple sclerosis (MS) lesions on T2-weighted FLAIR images [24].
@@
    Brain Mets (D2): Brain metastases imaged through T1, contrast enhanced gradient echo T1ce, contrast enhanced spin echo T1 and a T2 FLAIR sequence acquired in the Stanford University Hospital [11].
@@
    Hippocampus (D3): The Hippocampus dataset, task 4 of the medical segmentation decathlon (MSD) [1], contains delineations of the anterior and posterior Hippocampus in T1 weighted MRI [29].
@@
    Atlas22 (D4): Anatomical Tracings of Lesions After Stroke (ATLAS) on T1 weighted images. We use the Atlas R2.0 dataset from Liew et al. [22].
@@
    CrossModa (D5): Delineations of intra-meatal and extra-meatal vestibular schwannoma tumors and cochlea delineations in contrast-enhanced T1 weighted MRI [7].
Of all these datasets, we set aside a hold-out test set comprising 20% of all images before the start of the method development process. The remaining images were split 80/20 into training and validation sets for the development process.
@@
    Cosmos (D6): This dataset features carotid vessel wall segmentation and atherosclerosis diagnosis, of which we use the contours to evaluate segmentation performance [3].
@@
    HaNSeg (D7): This dataset contains segmentations of 30 organs at risk (OAR), with associated T1 MRI and CT, of which we only use the MR images for model development [26].
@@
    Isles22 (D8): This dataset contains annotations of ischemic stroke lesions, with associated diffusion-weighted imaging (DWI), apparent diffusion coefficient (ADC) and T2 FLAIR [16].
@@
    HNTS-MRG24 (D9): T2-weighted MR images of pre-treatment oropharyngeal cancer and metastatic lymph nodes with associated annotations [38].
@@
For test datasets D6-D12 we use an 80/20 split for fine-tuning and for testing. A validation set is omitted as we fine-tune each method only once without any interventions. For D13 all data is used for testing, as no training is conducted.
3 Related Work
In this section, we revisit currently established MAE pre-training paradigms. Then, we survey existing MAE pre-training paradigms in the 3D medical domain, before expanding this to a broader set of pre-training paradigms for 3D medical image segmentation.
3.1 Masked Autoencoders
Auto-encoders have been established for a long time through the early pioneering works of Vincent et al. [37] but recently garnered significant attention through He et al. [14], showing that a simple MAE pre-training paradigm is scalable and provides substantial performance improvements. Subsequently, follow-up works extended these findings to video data [9, 34] as well as language-image data [21]. These earlier successes primarily relied on Transformer architectures, as their sequence modeling offered significant VRAM savings. CNNs on the other hand are less suited to such a pre-training paradigm as masking disrupts the 2D data-structure they compute on. Despite this Woo et al. [41] and Tian et al. [32] were able to extend MAE pre-training successfully to CNNs. Woo et al. [41] introduced their fully convolutional masked autoencoder (FCMAE) framework, while Tian et al. [32] introduced adaptations to the convolution computation and normalization to make it resilient to masking.
3.2 Masked Autoencoder In 3D Medical
The earliest work using autoencoders for self-supervised pre-training in the medical domain was Models Genesis (MG) conducted by Zhou et al. [44] in 2019. They introduced reconstruction of in-painting, out-painting, and denoising of local intensity shifts as their pre-training task. Many approaches followed, with Swin UNETR by Tang et al. [31] proposing a joint in-painting, rotation, and contrastive objective for Swin Transformer pre-training. Chen et al. [5] proposed a basic masking and reconstruction strategy applied to 3D Vision Transformers (ViT) [8] termed (3DMAE) and Zhuang et al. [45] proposed to use a combination of masked global and local views in conjunction with a reconstruction and consistency loss to pre-train 3D ViTs, named GL-MAE. MAPSeg by Zhang et al. [43] combines a masked autoencoder objective and pseudo-labeling to address domain shifts in a federated setting. More recently, in concurrent work, Munk et al. [23] proposed pre-training CNNs with a vanilla masking and reconstruction objective without any adaptations on a large brain dataset (AMAES) and Tang et al. [30] presented a sparse MAE with reconstruction objective pre-trained on abdomen CT data (HySparK).
3.3 Other Pre-Training For 3D Medical
Additionally, self-supervised learning approaches for 3D medical images include Zhuang et al. [46], who introduced a training objective aimed at recovering the orientation and rotation of shuffled and rotated sub-volumes, inspired by a Rubik’s cube. Wang et al. [39] proposed a method that involves mixing two images at varying patch rates and predicting the mixing categories via a segmentation objective, termed Volume Fusion (VF). Geometric Visual Similarity Learning (GVSL) of He et al. [15] leveraged anatomical consistency by learning to register two images of the same field of view and using local similarity as a training objective. More recently, Wu et al. [42] suggested partitioning a volume into non-overlapping base-crops and generating additional overlapping sub-crops, with a training objective to predict the degree of overlap between sub-crops and base-crops (VoCo).
@@ 4 Revisiting 3D MAEs
(a) Sparsification: The introduction of all sparsification components showed best results, with most improvements from adding the additional densification conv layer between encoder and decoder.
@@
(b) Masking ratio: Masking ratios between 60% and 75% worked best when choosing static ratios. Overall, a dynamic range, including higher masking ratios, performs the best.
@@
(c) Development performance: When comparing our final S3D model against the baseline methods, trained equally, S3D exceeds all baselines on the development datasets.
SSL Method No (Dyn.) No (Fix.) S3D* S3D
Dataset Dice Similarity Coefficient (DSC)
MS FLAIR (D1) 57.81 59.82 59.75 60.35
Brain Mets (D2) 63.66 56.53 64.20 65.24
Hippocampus (D3) 89.18 89.24 89.45 89.60
Atlas22 (D4) 63.28 65.52 66.61 66.95
CrossModa (D5) 85.64 83.44 83.61 84.08
Cosmos22 (D6) 60.28 78.17 80.01 80.00
ISLES22 (D7) 77.94 79.44 78.94 79.70
Hanseg (D8) 59.00 61.85 61.27 62.11
HNTS-MRG24 (D9) 66.73 65.90 67.03 68.62
BRATS24 Africa (D10) 93.07 92.51 92.49 92.19
Avg. DSC 71.66 73.24 74.34 74.88
Avg. Rank 3.2 2.9 2.4 1.5
Masked autoencoders (MAEs) are a well-established pre-training paradigm in the natural imaging domain and in the medical image segmentation domain for transformers. In this section, we investigate this paradigm and optimize it for 3D medical image segmentation using a ResEnc U-Net architecture of Isensee et al. [19].
@@
MAEs are trained by masking an input image to a certain degree and training the network to reconstruct the occluded regions, minimizing deviations between reconstruction and the original image. In our experiments, we train the MAE with an L2-Loss in the z-score normalized voxel space and only calculate the reconstruction loss where regions were masked. Moreover, we do not remove skip-connections, following the general consensus of Woo et al. [41], Tian et al. [32], and He et al. [14]. The default hyperparameters (as used by the model denoted in gray in Tab. 2(c)) are learning rate 1e-2, weight decay 3e-5, batch size 6, SGD optimizer with Nesterov momentum 0.99, masking ratio of 75% trained with a PolyLR schedule for 250k steps (this represent 1000 epochs in the nnU-Net framework) and minor spatial augmentations of affine scaling, rotation and mirroring.
@@
When masking the input image, CNNs are not able to ignore the masked regions in the same manner as transformers can. To address this, Tian et al. [32] proposed to adapt the CNN architectures to better fit the sparse inputs: (a) Sparse Convolutions and Normalization:Through the receptive field of convolutions masked-out regions are iteratively eroded from their boundaries. By re-applying the masked regions after every convolution this problem can be resolved. Moreover, the masks can introduce a problematic shift in the normalization layer statistics due to the introduced zero values. To resolve this normalization is constrained to only consider the non-masked values. (b) Mask Token:Instead of feeding the light-weight decoder feature maps with zeroed mask regions, the regions are densified by filling them with a learnable mask token, simplifying the reconstruction task of the decoder. (c) Densification Convolution:After filling the masks with the Mask Token and before passing the feature maps to the decoder, a [3x3x3] convolution is applied to the feature maps at every resolution except the highest resolution to prepare the representations for decoding.
Results of these changes are visualized in Tab. 2(a). The adaptations are introduced iteratively, meaning the ’MaskToken’ ablation is only applied together with the Sparse Convolutions and Normalization. It can be observed that the full set of adaptations improves performance by an average of $0.3$ DSC points across our development datasets. Subsequently, all adaptations are kept and the following evaluations are presented with these changes applied.
@@
The masked region is determined by sampling randomly in the CNN’s bottleneck of shape [5x5x5] and up-sampling these regions to the input resolution, to ensure the masks align in the bottleneck of the CNN architectures. This results in masking regions of [32x32x32] voxels of non-overlapping regions in the input. As a sampling strategy, we follow random masking, as previous work showed no benefit of structured masking for images nor videos [14, 9]. In the scope of the development phase we explore 5 static masking ratios between 30 and 90%, and evaluate a dynamic masking ratio randomly masking between 60% and 90%.
Results are presented in Table Tab. 2(b) and highlight that the masking ratios of 60%, 75%, and the dynamic masking ratio of 60% to 90% perform equally well. Due to the highly similar performance, we choose to proceed with a dynamic masking ratio over the static masking ratio, due to expecting this masking to be more difficult to learn and with the upcoming scaling experiment in mind. We refer to this model as Spark3D (S3D) going forward.
Table 3: Fine-tuning matters. We compare various combinations of weight transfer and fine-tuning schedules. Transfer: [Uncaptioned image] Transfer all weights, [Uncaptioned image] Transfer encoder weights only. Warm-Up and Fine-tuning: [Uncaptioned image] Only decoder weights adapted during fine-tuning, [Uncaptioned image] Encoder and decoder weights adapted during fine-tuning. $\ddagger$: nnU-Net default (Dynamic planning.)
Transfer 1. Warm-Up 2. Warm-Up Fine-tuning Max. LR D1 D2 D3 D4 D5 Avg